OpenStudy (anonymous):

Problem 54/120 A committee of four is chosen randomly from a group of six males and eight females. What is the probability that the committee includes either all males or all females? (Round your answer to four decimal places.)

OpenStudy (kropot72):

The probability of all males is given by: \[\large P(4\ males)=\frac{6C4}{14C4}\] The probability of all females is given by: \[\large P(4\ females)=\frac{8C4}{14C4}\] The events '4 males' and '4 females' are mutually exclusive. Therefore P(4 males or 4 females) = P(4 males) + P(4 females)

OpenStudy (anonymous):

@kropot72 Thank you !!! your a life saver !!!

OpenStudy (kropot72):

You're welcome :)